OpenStudy (anonymous):

Change the percent to a decimal. 12 3/5% A.1.26 B.0.126

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

\[12\frac{3}{5}\% = \frac{(12\cdot 5)+3}{5}\%\]First convert that. what does it give you?

OpenStudy (anonymous):

my answer is 0.126 is it correct?

OpenStudy (jhannybean):

Don't worry about your answer, we can check that later. what do you get as your simplified fraction?

OpenStudy (jhannybean):

No.

OpenStudy (anonymous):

1260

OpenStudy (jhannybean):

What is 12 x 5

OpenStudy (anonymous):

60

OpenStudy (jhannybean):

+3?

OpenStudy (anonymous):

63

OpenStudy (jhannybean):

ok, so you have \[\frac{63}{5}\%\]When you write a percent as a fraction, you divide by 100

OpenStudy (jhannybean):

So what is 5 x 100?

OpenStudy (anonymous):

500

OpenStudy (jhannybean):

\[\frac{63}{5}\% = \frac{63}{500}\]

OpenStudy (jhannybean):

Just convert 63/500 as a decimal using your calculator, and you will have your answer.

OpenStudy (anonymous):

0.126